If tan(A/2) = 5/6, tan(B/2) = 20/37, prove that a+c = 2b 

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear

we know

tanA/2 ={(s-b)(s-c)/s(s-a)}1/2

tanB/2 ={(s-a)(s-c)/s(s-b)}1/2

now first calculate

 tanA/2 * tanB/2 = 5/6  * 20/37

61a -161c + 61b =0   .................1

 

now calcualte

 tanA/2 / tanB/2  = 5/6 ÷ 20/37

put the value and simplify

 61a -13c -61b =0  ....................2

 

 now solve 1 and 2

a/87  = b/74 =c/61

 

 calculate a+c = [87/74 + 61/74 ]b

                     =2b

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you the answer and detailed
solution very  quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin